OpenStudy (anonymous):

Use De Moivre's theorem to write the complex number in trigonometric form. (cos (4pi/7) + i sin (4pi/7)) ^5 A. ( cos (20pi/7) + i sin (20pi/7)) B. 5 ( cos (4pi/7) + i sin (4pi/7)) C. ( cos (4pi/35) + i sin (4pi/35)) D. 5 ( cos (20pi/7) + i sin (20pi/7))

OpenStudy (anonymous):

\[[r(\cos \theta+i \sin \theta)]^{n}\]